A washing machine selling for RM800 cash was purchased through an instalment plan in which

10% down payment was made. The buyer made 20 weekly payments of RM40 each to settle

the balance. Calculate

(a) the instalment price

(b) the total interest charged

(c) the simple interest rate charged

(d) the effective rate charged


1
Expert's answer
2021-01-04T19:32:00-0500

(a) The total installment price is the total amount you pay


"20\times RM40+0.1\times RM800=RM 880"

(b)

The total interest charged =RM880−RM800=RM80


(c)

The cash price "=RM800"

Cash down payment "=0.1\times RM800=RM80"

Balance to be paid "=RM800-RM80=RM720"

If the rate of interest per annum under instalment plan is "r\%," then


"720+720\times\dfrac{r}{100}\times\dfrac{20}{52}=20\times40"

 The simple interest rate charged

"r=28.89\%"

(d)

The Effective Annual Interest Rate "=(1+\dfrac{0.2889}{52})^{52}-1=0.3339"

The Effective Annual Interest Rate "=33.39\%"